OpenStudy (anonymous):

solve the equation 36x^2-4=0

hero (hero):

Hint: Difference of squares a^2 - b^2 = (a+b)(a-b)

OpenStudy (inkyvoyd):

add 4 to both sides. divide both sides by 36 simplify the fraction tak the plus minus square root of both sides

hero (hero):

In this case, we have 36x^2 - 4 = (6x + 2)(6x -2) = 0

OpenStudy (inkyvoyd):

Hero's method works as well, and is faster, but you need to have factoring patterns memorized.

hero (hero):

Both methods work equally well.

hero (hero):

I use them interchangably

OpenStudy (inkyvoyd):

Also, note that plus minus square root means you have two answers; many tend to forget one

OpenStudy (anonymous):

final answer x= 1/3, =1/3

hero (hero):

actually, I think x = 1/3, -1/3

OpenStudy (anonymous):

typo thats what i meant

OpenStudy (nastech):

36x^2-4=0 4(9x^2-1)=0 9x^2-1=0 9x^2=1 x^2=1/9 x=1/3

hero (hero):

Ya, I know @parksbap that's why I gave you a medal. But I was stating the fact just for the record. Not necessarily for you but for others who are also viewing the question and may be trying to solve it themselves.

OpenStudy (anonymous):

thanks

OpenStudy (inkyvoyd):

Also, hedgehog.

OpenStudy (nastech):

actuallyit is 1/3 or -1/3